### v2

@@ 5.1.5 Choice of values for Tuning Parameters
Figure 2: Value of $p_{effective}$ (the effective false positive rate overhead) in our correctness benchmark at various parameter settings given $N$=24,956 documents and $p=$1e-5
@@
It is possible that one could properly tune these deduplication methods further using incremental changes in their parameter settings. However, the expensive runtime of n-gram techniques on large corpora coupled with the need to comb over parameter settings in small increments points to the inherent difficulty and cost associated with properly tuning these methods for maximal deduplication performance. Meanwhile, MinHashLSH-based techniques readily admit parameter tuning with relatively low runtime and memory costs. Furthermore, MinHashLSH and LSHBloom both benefit from established theoretical analysis that effectively bounds their false positive and false negative rates given the values for the Jaccard similarity threshold and number of permutations. MinHashLSH-based techniques offer the greatest opportunity to tune parameters for the best deduplication performance, particularly when faced with a constrained parameter tuning budget. This is particularly true for LSHBloom, which is significantly cheaper to run than MinHashLSH in terms of wall clock time and disk usage—particularly as we increase the size of our dataset to several million or several billion documents (see Section 5.3.1).
@@
Figure 3: F1 score for LSH techniques as a function of the number of permutations (x-axis) and Jaccard similarity threshold (y axis).
5.2 Deduplication Correctness
5.2.1 LSH Methods
The parametrization of methods such as MinHashLSH and LSHBloom has dramatic effects for their accuracy so tuning the parameters is a critical aspect of deploying them for large-scale deduplication efforts. Figures 3(a) and 3(b) show the F1 score for MinHashLSH and LSHBloom, respectively, with different settings for the Jaccard similarity threshold, $T$, and the number of permutation functions for MinHashing on the 25K dataset. As the number of permutation functions increases, the MinHash estimate of the Jaccard similarity becomes more accurate. However, there is a clear saturation point at 128 permutation functions, after which we observe only marginal performance increases. This is particularly relevant for LSHBloom, for which increasing the number of permutation functions blindly can lead to slight degradations in performance—due to the introduction of more bands in the signature matrix by the LSH function, requiring more Bloom filters, and therefore leading to an increase in the false positive rate. This can be avoided, however, by jointly tuning the false positive rate and the number of permutation functions. The threshold parameter affects the F1 score the most, as it defines the sensitivity of the deduplication algorithm to overlap in text content. For our particular benchmark, a threshold of 0.6 yields the best performance, as more stringent thresholds lead to a diminishing F1 score.
@@ 5.2.1 LSH Methods
Figure 4: F1 Score for N-Gram techniques as a function of N-Gram size (x axis) and overlap threshold (y axis).
5.2.2 Paragraph Level and N-Gram Methods
@@
Figure 5 shows the impact of adjusting the overlap threshold parameter for paragraph-level methods, DOLMA and CCNet, on the F1 score. Distinguishing documents at the granularity of individual paragraphs is highly error-prone and results in poor performance. N-gram methods such as DOLMA-ngram and CCNet do comparably better in terms of F1 score but still lag just behind the scores of the MinHashLSH-based techniques. Figures 4(a) and 4 show the results of varying both the n-gram size and overlap threshold parameters for each n-gram method. While n-gram methods enable a finer-grained comparison between individual documents and the rest of the corpus, these methods evaluate overlap by determining the portion of n-grams in a document that is exactly duplicated in the corpus. This is similar to the notion of evaluating the set overlap between documents but can be sensitive to the relative frequency of n-grams within a document, resulting in increased error. Moreover, both paragraph and n-gram methods demonstrate a lack of sensitivity to changes in their parameter settings in this benchmark, indicating bias in their predictions. Specifically, we find that paragraph-level methods tend to be biased toward marking documents as duplicates. In contrast, n-gram methods yield many false negatives at high thresholds and yield many false positives at low thresholds.
Figure 5: F1 score vs. threshold for paragraph-level techniques.
5.3 Deduplication Resource Usage
@@
Figure 6: Pareto plots for F1 score vs. resource usage.
As LLM training datasets grow ever larger, it is important that text deduplication strategies can scale effectively to larger quantities of documents. Here we work with the peS2o dataset Soldaini & Lo (2023) of roughly 8M full-text and 31M abstracts. We evaluate each deduplication method on subsets of peS2o of various sizes and compare them in terms of wall clock time and disk usage (e.g., for storing an index or Bloom filter). We do not evaluate the n-gram techniques for scaling capacity on peS2o due to their prohibitive cost. DataComp-LM, for example, takes one hour just to deduplicate 25,000 examples in our previous benchmark in Section 5.1.4. Even processing just 10% of peS2o using this technique would require 160 hours by extrapolation. We discuss the results for other methods below.
Figure 7: Disk size of index for different methods, as measured for peS2o subsets of increasing size.
Figure 8: Wall clock times for different methods, as measured for peS2o subsets of increasing size.
5.3.1 Scaling of Resource Usage
In this section, we study how resource usage grows with the number of documents. In these experiments, we set any threshold parameters to 0.8, the number of random permutations for our LSH techniques to 128, and $p$ = 1e-5 for all Bloom filters. Ideally, we would have informed these parameter settings based on the optimal settings from Section 5.1.4; however, scaling experiments on peS2o regularly took over a day to complete, limiting our ability to tune parameters and necessitating heuristic choices for parameter settings as we evaluated deduplication quality in parallel.
Figure 8 shows the wall clock time to deduplicate various subsets of peS2o with each method. Notably, MinHashLSH incurs the largest runtime while paragraph-level techniques such as DOLMA and CCNet are relatively fast. However, DOLMA and CCNet achieve this speed with significantly lower deduplication performance: see Section 5.1.4. Meanwhile, LSHBloom, which has comparable deduplication performance to MinHashLSH, is much faster. To deduplicate all of peS2o, LSHBloom takes just five hours, whereas MinHashLSH takes 14 hours—almost a 3$\times$ speedup. Figure 1 shows the breakdown of this time for MinHashLSH and LSHBloom; we can see that, as expected, the time to compute MinHashes is the same; however, the insertion and query latencies are significantly reduced in the Bloom filter-based index compared to the traditional LSHIndex.
Figure 7 shows the disk usage for each method. We again see a similar trend. The paragraph-level techniques maintain only a single Bloom filter which leads to extremely low disk usage. Meanwhile, MinHashLSH’s index grows steadily in size as the number of documents increases. LSHBloom’s index also grows linearly in the number of documents but at a reduced rate. When processing all of peS2o, LSHBloom requires just 1 GB vs. the 180 GB taken by MinHashLSH—a 180$\times$ reduction.
5.3.2 Estimated Resource Usage at Scale
Figure 9: Predicted wall clock times vs. number of documents (in Billions) for different methods
@@ 5.3.2 Estimated Resource Usage at Scale
Table 1: Predicted index sizes for 5B documents, for LSH and LSHBloom with different false positive overheads.
Technique Bloom filter false Index disk
positive overhead usage (GB)
MinHashLSH 23,008.04
LSHBloom 1e-5 160.51
LSHBloom 1e-10 295.30
LSHBloom 1e-15 425.35
Table 2: Predicted index sizes for 100B documents, for LSH and LSHBloom with different false positive overheads.
@@
MinHashLSH 460.16
LSHBloom 1e-5 3.21
LSHBloom 1e-10 5.91
LSHBloom 1e-15 8.51
5.4 Performance vs. Quality Tradeoffs at Scale
Here we investigate tradeoffs between the costs of various deduplication techniques (in terms of runtime and disk usage) and their deduplication performance. Figures 6(a) and 6(b) plot the maximum F1 score for each deduplication technique against runtime and disk usage, respectively. MinHashLSH and LSHBloom dominate the alternative techniques, achieving the highest F1 scores while also exhibiting competitive runtime. Additionally, LSHBloom displays a comparable F1 score to traditional MinHashLSH while using just a fraction of the disk space for its index. As shown in Section 5.3.1, this is a crucial distinction, as rapidly increasing disk usage renders deduplication efforts infeasible at the scale of many millions or billions of documents.
6 Conclusions
We have introduced a new text deduplication technique, LSHBloom, that offers significant runtime and space savings over the current state-of-the-art technique, MinHashLSH. We also conducted a thorough study of deduplication quality between our technique and other state-of-the-art deduplication algorithms commonly used in LLM training pipelines, using a custom benchmark of 25,000 scientific articles. We find that even with extensive parameter tuning for conventional MinHashLSH, LSHBloom is extremely competitive in terms of its deduplication quality, with an F1 score within 10% of the best achieved by MinHashLSH, while using far less storage and runtime. Using the peS2o dataset, we find that LSHBloom can feasibly scale to deduplicating billions of documents due to its low space requirements (a 180$\times$ space reduction from MinHashLSH) and relatively low wall clock time (3$\times$ speedup over MinHashLSH). In contrast, alternatives demonstrate either poor deduplication quality, excessive runtime, or prohibitive space requirements for extreme-scale datasets of this size. In the future, we aim to further reduce the wall clock time of LSHBloom by employing parallelization over subsets of text datasets when inserting them into our index—an approach that we expect to enable significant acceleration.

### v3

@@ 5.1.5. Choice of values for Tuning Parameters
Figure 2. Value of $p_{effective}$ (the effective false positive rate overhead) in our correctness benchmark at various parameter settings given NN=24,956 documents and pp=1e-5
@@
It is possible that one could properly tune these deduplication methods further using more incremental changes in their parameter settings. However, the expensive runtime of n-gram techniques on large corpora coupled with the need to comb over parameter settings in small increments points to the inherent difficulty and cost associated with properly tuning these methods for maximal deduplication performance. Meanwhile, MinHashLSH-based techniques readily admit parameter tuning with relatively low runtime and memory costs. Furthermore, MinHashLSH and LSHBloom both benefit from established theoretical analysis that effectively bounds their false positive and false negative rates given the values for the Jaccard similarity threshold and number of permutations. MinHashLSH-based techniques offer the greatest opportunity to tune parameters for the best deduplication performance, particularly when faced with a constrained parameter tuning budget. This is particularly true for LSHBloom, which is significantly cheaper to run than MinHashLSH in terms of wall clock time and disk usage—particularly as we increase the size of our dataset to several million or several billion documents (see subsubsection 5.4.1).
@@
Figure 3. F1 score for LSH techniques as a function of the number of permutations (x-axis) and Jaccard similarity threshold (y axis).
5.2. Hyperparameter Tuning Results (24K Dataset)
5.2.1. LSH Methods
The parametrization of methods such as MinHashLSH and LSHBloom has dramatic effects for their accuracy so tuning the parameters is a critical aspect of deploying them for large-scale deduplication efforts. Figures 3(a) and 3(b) show the F1 score for MinHashLSH and LSHBloom, respectively, with different settings for the Jaccard similarity threshold, $T$, and the number of permutation functions for MinHashing on the 24K dataset. As the number of permutation functions increases, the MinHash estimate of the Jaccard similarity becomes more accurate, and performance generally improves, which is consistent with the theory of MinHashing (broder1998min). Correspondingly, we use 256 MinHash permutations in our subsequent experiments. The threshold parameter affects the F1 score the most, as it defines the sensitivity of the deduplication algorithm to overlap in text content. For our particular benchmark, a threshold of 0.5 yields the best performance; more stringent thresholds lead to a diminishing F1 score.
@@ 5.2.1. LSH Methods
Figure 4. F1 Score for N-Gram techniques as a function of N-Gram size (x axis) and overlap threshold (y axis).
5.2.2. Paragraph Level and N-Gram Methods
@@
Figure 5 shows the impact of adjusting the overlap threshold parameter for paragraph-level methods, DOLMA and CCNet, on the F1 score. Distinguishing documents at the granularity of individual paragraphs is highly error-prone and results in poor performance. The best F1 for paragraph-level methods is achieved by setting a low threshold of 0.2. N-gram methods such as DCLM do comparably better in terms of F1 score but still lag just behind the scores of the MinHashLSH-based techniques. Interestingly, DOLMA and DOLMA-ngram perform very similarly. Figures 4(a) and 4 show the results of varying both the n-gram size and overlap threshold parameters for each n-gram method. While n-gram methods enable a finer-grained comparison between individual documents and the rest of the corpus, these methods evaluate overlap by determining the portion of n-grams in a document that is exactly duplicated in the corpus. This is similar to the notion of evaluating the set overlap between documents, but can be sensitive to the relative frequency of n-grams within a document, resulting in increased error. Moreover, paragraph-level techniques demonstrate a lack of sensitivity to changes in their parameter settings in this benchmark, indicating bias in their predictions. Specifically, we find that paragraph-level methods tend to be biased toward marking documents as duplicates.
Table 1 shows the parameter settings for each technique that we use for the remainder of our experiments.
Figure 5. F1 score vs. threshold for paragraph-level techniques.
Table 1. Best Settings for Each Deduplication Technique
Technique N-gram Size Threshold
MinHashLSH 1 0.5
LSHBloom 1 0.5
Dolma-Ngram 5 0.2
DCLM 5 0.2
Dolma - 0.2
CCNet - 0.2
5.3. Deduplication Correctness (50K Dataset)
(a) Precision
(b) Recall
(c) F1
Figure 6. Correctness metrics for deduplication techniques for various levels of duplication.
5.3.1. Deduplication Quality
Here we evaluate the correctness of each technique using metrics such as their precision, recall, and F1 score. We do so by constructing datasets of 50,000 documents each with varying levels of duplication between 10% and 90%. We construct these datasets using the procedure described in subsubsection 5.1.4.
Figure 6 illustrates the precision, recall, and F1 score of each technique as we vary the proportion of duplicate documents in our dataset. Notably, in terms of F1 score (see 6(c)), both MinHashLSH and LSHBloom outperform all other methods—only being slightly outperformed by DCLM and Dolma-Ngram (whose performance tracks closely together) when the vast majority ($60\%$) of the dataset consists of duplicates, which is highly unlikely in real-world scenarios. MinHashLSH and LSHBloom also achieve nearly identical F1 scores across experimental settings, exemplifying that our technique can closely match the performance of MinHashLSH while affording considerable gains in scalability (to be shown below in subsection 5.4).
Looking more closely, 6(a) illustrates that MinHashLSH and LSHBloom tend to consistently outperform the other techniques in terms of precision, which is particularly noteworthy for those settings where duplicates are rare, such as when duplicates only comprise 10 or 20% of the dataset. The next best technique in terms of precision is Dolma’s paragraph-level deduplication technique, with its precision often tracking closely with that of the MinHash-based techniques. However, Dolma’s dismal recall scores (<0.4<0.4) suggest that this is likely the result of under-identifying duplicate instances rather than being a reflection of greater capacity for discernment. DCLM, Dolma-Ngram, and CCNet tend to underperform the other techniques in terms of precision. Note that there is a natural tendency for precision scores to increase as the proportion of duplicates in the dataset increases. This is due to a greater number of true positives being present in the dataset.
In terms of recall, 6(b) demonstrates that DCLM and Dolma-Ngram tend to perform the best—though notably their relatively stable recall scores come at the cost of decreased precision, particularly when duplicates are rare. This indicates that, while these techniques identify most duplicates, they struggle to differentiate between duplicate and non-duplicate entities. Meanwhile, MinHashLSH and LSHBloom provide virtually identical recall scores to one another between 0.7 and 0.9 while simultaneously demonstrating strong precision. Paragraph-level techniques demonstrate lower recall than all other techniques by a wide margin.
Our experimental results indicate that LSHBloom achieves deduplication performance comparable to the state-of-the-art—indeed, LSHBloom achieves essentially identical performance to MinHashLSH, one of the most popular choices for text deduplication, with only a marginal degradation in precision.
5.3.2. Performance vs. Quality Tradeoffs
Here we investigate tradeoffs between the costs of various deduplication techniques (in terms of runtime and disk usage) and their deduplication performance. We visualize data from our evaluation of the balanced dataset of 50,000 documents.
Figures 7(a) and 7(b) plot the F1 score for each deduplication technique against runtime and disk usage, respectively. MinHashLSH and LSHBloom dominate the alternative techniques, achieving the highest F1 scores while also exhibiting competitive runtime. LSHBloom runs faster than MinHashLSH and this runtime advantage becomes more exaggerated at scale (see subsection 5.4). LSHBloom also displays a comparable F1 score to traditional MinHashLSH while using just a fraction of the disk space for its index. As shown below in subsection 5.4, this is a crucial distinction, as rapidly increasing disk usage renders deduplication efforts infeasible at the scale of many millions or billions of documents.
Our results indicate that LSHBloom intelligently discerns between duplicate and non-duplicate documents (as demonstrated by its high precision scores) and can adeptly identify the majority of duplicate documents (as demonstrated by its strong recall scores). Our results also illustrate that the main difference in performance between MinHashLSH and LSHBloom comes in the form of slightly reduced precision for the latter. Intuitively, this makes sense as LSHBloom yields slightly more false positive predictions due to its use of the Bloom filter. Overall, however, both yield similar F1 scores. This, combined with LSHBloom’s greatly reduced resource usage compared to MinHashLSH, positions it as the best choice for text deduplication use-cases.
5.4. Deduplication Resource Usage at Scale
@@
Figure 7. Pareto plots for F1 score vs. resource usage.
As LLM training datasets grow ever larger, it is important that text deduplication strategies can scale effectively to larger quantities of documents. Here we work with the peS2o dataset (peS2o) of roughly 39M documents. We evaluate each deduplication method on subsets of peS2o of various sizes and compare them in terms of wall clock time and disk usage (e.g., for storing an index or Bloom filter). We do not evaluate the n-gram techniques for scaling capacity on peS2o due to their prohibitive cost. DataComp-LM, for example, requires almost three hours just to deduplicate 50,000 examples in our previous benchmark in subsection 5.3. Even processing just 10% of peS2o using this technique would require over 200 hours by extrapolation. We discuss the results for other methods below.
Figure 8. Disk size of index for different methods, as measured for peS2o subsets of increasing size.
Figure 9. Wall clock times for different methods, as measured for peS2o subsets of increasing size.
5.4.1. Scaling of Resource Usage
In this section, we study how resource usage grows with the number of documents. In these experiments, we use the same settings shown in Table 1. We set the effective false positive rate for all bloom filters to 1e-10.
@@ 5.4.1. Scaling of Resource Usage
[…]
5.4.2. Estimated Resource Usage at Scale
Figure 10. Predicted wall clock times vs. number of documents (in Billions) for different methods
[…]
Table 2. Predicted index sizes for 5B documents, for LSH and LSHBloom with different false positive overheads.
@@ 5.4.2. Estimated Resource Usage at Scale
MinHashLSH 277.68
LSHBloom 1e-5 8.33
LSHBloom 1e-8 12.11
LSHBloom 1/5B 15.50
Table 3. Predicted index sizes for 100B documents, for LSH and LSHBloom with different false positive overheads.
Technique Bloom filter false Index disk
positive overhead usage (TB)
MinHashLSH 555.35
LSHBloom 1e-5 16.66
LSHBloom 1e-8 24.21
LSHBloom 1/100B 31.76
6. Conclusions
[…]
In the future, we aim to further reduce the wall clock time of LSHBloom by employing parallelization over subsets of text datasets when inserting them into our index—an approach that we expect to enable significant acceleration. Additionally, though very useful, our synthetic benchmark has drawbacks—namely, in its biased and limited notion of what constitutes a proper duplicate. Ideally, future work will involve benchmark datasets that cover more diverse forms of duplication in proportion to their representation ”in the wild”—that is, in real-world LLM pre-training data pipelines.
Appendix A Comparison of Data Deduplication Pipelines for LLM Pre-training
Table 4. Comparative Analysis of Data Deduplication Pipelines for LLM Pre-training
Dataset / Model Primary Data Sources Key Cleaning & Filtering Steps Dedup. Level Deduplication Technique(s) Key Parameters & Implementation Notes Citations
C4 (for T5) Common Crawl (April 2019 snapshot) Heuristic filtering: remove lines without terminal punctuation, lines << 3 words, docs << 5 sentences, docs with ”Lorem ipsum”, and docs containing words from a blocklist. Language filtering for English (>>99% probability). None Heuristic Filtering Only No explicit deduplication algorithms were used. The focus was on rule-based filtering. They remove all but one instance of any three-sentence span that occurs more than once in the dataset. (t5_c4_raffel2020)
The Pile (for GPT-Neo) 22 diverse sources including Common Crawl (Pile-CC), Books3, arXiv, GitHub, OpenWebText2, and PubMed Central. Source-specific cleaning (e.g., converting TeX to Markdown for arXiv, parsing top answers for StackExchange). The entire dataset was filtered for duplicates. Document MinHashLSH The OpenWebText2 component specifically used in-memory MinHashLSH via the datasketch library for near-duplicate detection. (gao2020pile800gbdatasetdiverse)
GPT-3 A mix of Common Crawl (filtered), WebText2, Books1, Books2, and Wikipedia. Quality filtering was applied to the Common Crawl portion. Document MinHashLSH Spark’s MinHashLSH implementation with 10 hashes. (brown2020language)
RedPajama / SlimPajama (LLaMA Replication) LLaMA sources: Common Crawl, C4, GitHub, Books, arXiv, StackExchange. Multi-stage pipeline: NFC Unicode normalization, filtering documents << 200 characters. Document Exact Hashing (Bloom Filter) & Fuzzy Deduplication (MinHashLSH) RedPajama-V2 uses a Bloom filter for efficient exact deduplication in addition to MinHashLSH. SlimPajama uses a distributed MinHashLSH implementation with 13-grams. (together2023redpajama; cerebras2023slimpajama)
CCNet CommonCrawl Language identification (for splitting into monolingual datasets) using lightweight classifiers. LM-based quality filtering at the document level. Exact (SHA1 Hashing) Line-level Computes SHA1 hashes for each line of text and performs exact deduplication at the line-level using these hash values. (wenzek2020ccnet)
Llama 3 Constructed ”from a variety of data sources containing knowledge until the end of 2023”. PII and safety filtering. Heuristic filtering: removed lines with repeated content, removed lines with ”dirty” words, filtered out documents with excessive numbers of outlier tokens compared to the rest of the training corpus. Document and Line-Level URL (metadata), MinHash, CCNet The paper mentions the use of URL deduplication, MinHash for document-level deduplication, and CCNet for line-level deduplication. (dubey_llama_2024)
Dolma Collected from CommonCrawl, Github, Reddit, Wikipedia, among other sources. Remove data containing PII or toxic material. Not Specified Paragraph Deduplicates paragraphs via exact-matching (Bloom filter). Alternatively, deduplicates paragraphs by measuring their n-gram overlap with the rest of the corpus, also using a Bloom filter. Preprocess some web data using CCNet and metadata-based deduplication. (soldaini_dolma_2024)
DataComp-LM (DCLM) Common Crawl (provides a candidate pool of 240T+ tokens). A benchmark for data curation strategies. The DCLM-Baseline dataset was created using model-based filtering (data quality classifiers) to select high-quality data. Paragraph and Document-level Baseline dataset is deduplicated using Bloom filters to measure n-gram overlap of a paragraph/document with the rest of the dataset. The DCLM-Baseline dataset of ~4T tokens is combined with StarCoder and ProofPile2. (li2024datacomp)
FineWeb / FineWeb-Edu / FineWeb2 Common Crawl (96 snapshots for FineWeb, ~100 for FineWeb2). URL filtering (blocklist), language classification (fastText), and quality/repetition filters. FineWeb-Edu uses a classifier trained on Llama-3-70B annotations. Document Fuzzy Deduplication (MinHash) Extensive ablation studies. The final strategy involved deduplicating each Common Crawl snapshot *individually*, as global deduplication was found to discard high-quality data from older snapshots. (penedo2025fineweb2pipelinescale)
Lee et al. (2021) C4, RealNews, LM1B, Wiki-40B-en N/A (focus was on the deduplication technique itself). Substring (Exact) Suffix Array Removed all but one occurrence of any 50-gram (sequence of 50 tokens) that appeared in multiple documents. This method targets boilerplate text. (lee2021deduplicating)
D4 (SemDeDup) Common Crawl (CC-dedup) An initial pass of MinHash-based deduplication was applied before the main technique. Document (Semantic) Semantic Deduplication (Embeddings + K-Means) A multi-stage method that first applies semantic deduplication (SemDeDup) to remove dense clusters of semantically redundant information. (tirumala2023d4)
LSHBloom peS2o (a dataset of 39 million technical documents) N/A (focus was on the deduplication technique itself). Document (Approximate) MinHashLSH with Bloom Filter Index An extension of MinHashLSH that replaces the large, memory-intensive LSH index with a series of lightweight Bloom filters (one for each band). $12\times$ faster, $18\times$ less space usage.